# CKD: Contrastive Knowledge Distillation from A Sample-wise Perspective

### v1

@@ III-B3 Contrastive Formulation
For sample-wise triples, $\exists j,f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)\sim f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, we can obtain $\exists j,\rho_{j}\sim 0$ and a larger $B\times e^{-\rho_{j}}$ is achieved, when we sample negative samples that are from the same class. Our method is unnecessary for comprehensive negative space sampling with large batch sizes. ∎
Fig. 6 visualizes traditional class-wise and the proposed sample-wise formulations. Specifically, CRD [70] is a representative class-wise contrastive learning approach by using a memory bank. We can observe that class-wise formulations including CRD generate negative pairs from different classes. These negative samples often have low similarity due to varied semantic content, necessitating a larger batch size to increase $g_{i}$. In contrast, the sample-wise formulation selects negative samples from the same class, thereby exhibiting stronger relationships. Thus, our method does not require a large batch size. In fact, a large batch size can hinder our method by influencing intra-sample alignment with a large $\beta$. However, a small batch size may also lead to insufficient gradient magnitude.
@@
Specifically, given the logit space $\boldsymbol{T}\in\mathbb{R}^{c\times n}$, the negative logit space for the $i$-th sample is denoted as $\boldsymbol{T}_{/i}\in\mathbb{R}^{c\times(n-1)}$. Assume that there are $c$ classes, each containing $m$ samples. We can derive:
@@
Eq. 20 suggests that in the proposed method, the number of negative pairs sampled should heavily depend on the class number $c$ rather than the sample number $n$. Taking category correlations into consideration, the actual number of negative pairs sampled should be even lower than $c$.
Furthermore, compared to class-wise contrastive learning approaches, our method can be efficiently trained with well-designed triples. For classic contrastive learning methods with $n$ samples, a total of $\mathcal{O}(nm)$ triples should be selected for each sample:
$\binom{m-1}{1}\binom{n-m}{1}=nm-m^{2}-n+m\propto\mathcal{O}(nm)$ (21)
Eq. 21 shows that for each sample, we randomly select one sample within the same class including $m-1$ selections, and one sample from a different class, yielding a total of $\mathcal{O}(nm)$ triples. As for CRD, we only need to select one sample from different classes because the positive pair is obtained by using the same sample:
$\binom{1}{1}\binom{n-m}{1}=n-m\propto\mathcal{O}(n).$ (22)
However, the negative space $\mathcal{O}(n)$ is vast, necessitating the use of a memory bank with a size of $K$. As a result, $K$ triples are available for training each sample. In contrast, the proposed method requires sampling only $\mathcal{O}(c)$ triples:
$\binom{1}{1}\binom{c}{1}=c\propto\mathcal{O}(1),$ (23)
where $c\ll n$ We randomly shuffle data samples and select non-overlapping $c$ samples in order. Therefore, our sample-wise formulation exhibits efficient trainability in comparison to the class-wise formulation in contrastive knowledge distillation.
To sum up, we encapsulate the training procedure in Algorithm 1.
TABLE II: Classification accuracy (%) on the CIFAR-100 validation set for different teacher-student network architecture. $\mathrm{\Delta}$ denotes the improvement in performance compared to vanilla KD. Results are averaged over five trials.
Methods Teacher ResNet32$\times$4 WRN-40-2 VGG13 ResNet50 ResNet32$\times$4
79.42 75.61 74.64 79.34 79.42
Student ShuffleNetV1 ShuffleNetV1 MobileNetV2 MobileNetV2 ShuffleNetV2
70.50 70.50 64.60 64.60 71.82
Features FitNet[59] 73.59 73.73 64.14 63.16 73.54
AT[31] 71.73 73.32 59.40 58.58 72.73
RKD[55] 72.28 72.21 64.52 64.43 73.21
CRD[70] 75.11 76.05 69.73 69.11 75.65
OFD[24] 75.98 75.85 69.48 69.04 76.82
ReviewKD[7] 77.45 77.14 70.37 69.89 77.78
Logits Vanilla KD[25] 74.07 74.83 67.37 67.35 74.45
TAKD[51] 74.53 75.34 67.91 68.02 74.82
DKD[90] 76.45 76.70 69.71 70.35 77.07
CKD 77.72 77.47 70.11 70.62 78.18
$\mathrm{\Delta}$ +3.65 +2.64 +2.74 +3.27 +3.73
IV Experiments
IV-A Image Classification
IV-A1 Datasets
We first conduct experiments on CIFAR-100 [32] and ImageNet-1K [13]. The CIFAR-100 dataset contains 100 classes, each with 600 images. The training and validation sets comprise 50K and 10K images. The ImageNet-1K dataset consists of 1000 classes. The training set has about 1.28 million images, and the validation set contains 50k images.
TABLE III: Top-1 and Top-5 accuracy (%) on the ImageNet-1K validation set. ResNet-34 serves as the teacher, while ResNet-18 acts as the student. $\mathrm{\Delta}$ denotes the performance improvement compared to vanilla KD. All results are the average over three trials.
Baselines Features Logits
Teacher Student AT[85] OFD[24] CRD[70] ReviewKD[7] Vanilla KD[25] DKD[90] CKD $\mathrm{\Delta}$
Top-1 73.31 69.75 70.69 70.81 71.17 71.61 70.66 71.70 72.24 +1.58
Top-5 91.42 89.07 90.01 89.98 90.13 90.51 89.88 90.41 90.81 +0.93
TABLE IV: Top-1 and Top-5 accuracy (%) on the ImageNet-1K validation set. ResNet-34 serves as the teacher, while ResNet-18 acts as the student. $\mathrm{\Delta}$ denotes the improved performance compared to vanilla KD. All results are averaged over three trials.
Baselines Features Logits
Teacher Student AT[85] OFD[24] CRD[70] ReviewKD[7] Vanilla KD[25] DKD[90] CKD $\mathrm{\Delta}$
Top-1 76.16 68.87 69.56 71.25 71.37 72.56 68.58 72.05 72.97 +4.39
Top-5 92.86 88.76 89.33 90.34 90.41 91.00 88.98 91.05 91.36 +2.38
TABLE V: Object detection performance on the MS-COCO dataset. We employ the two-stage Faster RCNN with FPN as the detector. Teacher-student pairs are ResNet101 & ResNet-18, ResNet101 & ResNet50, and ResNet50 & MobileNetV2, respectively. $\mathrm{\Delta}$ represents the improved performance compared to ReviewKD.
Methods ResNet101 & ResNet18 ResNet101 & ResNet50 ResNet50 & MobileNetV2
@@ IV-A2 Experimental Setup
We evaluate eleven isomorphic or heterogeneous architectures on CIFAR-100 and compare with vanilla KD [25], AT [85], FitNet [59], RKD [55], CRD [70], OFD [24], ReviewKD [7], TAKD[51], and DKD [90].
For ImageNet-1K, we evaluate representatives, including ResNet34 & ResNet18, and ResNet50 & MobileNetV2. We provide comparisons with vanilla KD [25], AT [85], OFD [24], CRD [70], ReviewKD [7], and DKD [90]. Top-1 and top-5 recognition rates are reported.
@@
IV-A3 Implementation Details
All experiments are conducted on NVIDIA RTX 3090. For CIFAR-100, we set $\alpha$ to 100 and employ cosine annealing for learning rates. For homogeneous and heterogeneous networks, we set batch sizes to 64 and 32, with an initial learning rate of 0.05 and 0.01, respectively. For ImageNet-1K, we also use cosine annealing, with an initial learning rate of 0.2. We set $\alpha$ to 10 and a batch size to 512.
Figure 9: Comparison of training time per batch and accuracy on CIFAR-100. The teacher model is ResNet32$\times$4 and the student model is ResNet8$\times$4. Additional parameters are denoted after $+$, which are added based on the student’s foundation.
@@ IV-A3 Implementation Details
Methods Homogeneous architectures Heterogeneous architectures
Teacher ResNet56 WRN-40-2 WRN-40-2 VGG13 ResNet50 ResNet32$\times$4
@@
Student ResNet20 WRN-16-2 WRN-40-1 MobileNetV2 MobileNetV2 ShuffleNetV2
@@
TABLE VII: Ablation studies of triple design on the CIFAR-100 dataset. Experiments are conducted on both homogeneous and heterogeneous teacher-student pairs. Due to the fixed teacher, triples $\left(\mathbf{t}_{i},\mathbf{s}_{i},\mathbf{t}_{j}\right)$ is meaningless and excluded from this table.
Triples Homogeneous architectures Heterogeneous architectures
Teacher ResNet56 WRN-40-2 WRN-40-2 VGG13 ResNet50 ResNet32$\times$4
@@
Student ResNet20 WRN-16-2 WRN-40-1 MobileNetV2 MobileNetV2 ShuffleNetV2
@@
CKD$\left(\mathbf{s}_{i},\mathbf{t}_{i},\mathbf{s}_{j}\right)$ 71.76 75.31 74.40 70.04 68.93 78.02
CKD$\left(\mathbf{s}_{i},\mathbf{t}_{i},\mathbf{t}_{j}\right)$ 71.69 75.88 74.43 70.09 69.68 77.06
CKD$\left(\mathbf{t}_{i},\mathbf{s}_{i},\mathbf{s}_{j}\right)$ 72.12 76.28 75.14 70.11 70.62 78.18
@@ IV-A4 Performance on CIFAR-100
We conduct classification experiments by using eleven combinations of isomorphic or heterogeneous teacher and student network architectures, including ResNet [22], WideResNet [73], VGG [63], ShuffleNet [88, 47], and MobileNet [60].
Table I presents classification performance by using teacher and student architectures of the same style. We can observe that our method achieves competitive performance against feature-based and logit-based methods. Specifically, the proposed CKD improves vanilla KD accuracy by 1.36%, with a maximum gain of 3.43%. Moreover, CKD also outperforms CRD on the CIFAR-100 dataset, with an improvement of almost 1.00%.
Table II shows the classification performance of teacher-student networks with different styles. We see that our CKD method consistently yields improvements across teacher-student network pairs compared to vanilla KD. Specifically, for different architectures, CKD achieves an average improvement of 3.21% over vanilla KD, with an impressive maximum gain of 3.73%.

### v2

@@ III-B3 Contrastive Formulation
For sample-wise triples where $\exists j,f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)\sim f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, we can obtain $\rho_{j}\sim 0$ and a larger $B\times e^{-\rho_{j}}$ is achieved, when sampling negative samples from the same class. Our method eliminates the need for comprehensive negative space sampling with large batch sizes. This finding aligns with the theoretical insights established by Robinson et al.[60], who demonstrated that difficult negative samples that are close to the anchor can provide more information and accelerate the gradient update. ∎
Some works have revealed that each image is distinctive and can vary significantly from other images within the same semantic class [80, 48]. It suggests that sample-wise contrastive learning could potentially outperform class-wise approaches by preserving intrinsic structural information in this case. Specifically, class-wise methods, which enforce uniform aggregation of intra-class samples toward a centroid, risk distorting the natural semantic structures inherent in the data.
Other research suggests that negative samples in the same class would harm contrastive learning, as the distance between same-class samples may become excessively large [28]. Our method addresses this challenge through teacher-guided structural constraints. When feature representations $\boldsymbol{s}_{i}$ and $\boldsymbol{s}_{j}$ from the same class move away from each other, their respective convergence towards corresponding teacher anchors $\boldsymbol{t}_{i}$ and $\boldsymbol{t}_{j}$ effectively constrains their relative positions. Consequently, this process leads to the alignment between the student’s intra-class semantic structure and that of the teacher.
Fig. 6 illustrates the differences between class-wise and sample-wise contrastive paradigms. Traditional class-wise approaches like CRD [72] employ memory banks to construct negative pairs through class exclusion, i.e., $\mathbf{x}_{j}\in\mathcal{C}_{k}$ where $k\neq y_{i}$ and $\mathcal{C}_{k}$ is $k$-th category of samples with
$\mathbb{E}\left[f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)\right]\rightarrow 0\quad\forall j\in\mathcal{C}_{k\neq y_{i}}.$ (20)
The class-wise formulation generates negative pairs from different classes, and these negative samples often exhibit low similarity due to diverse semantic content, forcing gradient enhancement through batch size scaling $g_{i}\propto B$. In contrast, the sample-wise formulation may sample negative pairs from the same class as,
$\mathbb{E}\left[f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)\right]\gg 0.$ (21)
This non-zero similarity naturally amplifies gradient signals without batch size dependency. Therefore, our method does not require a large batch size. In fact, the proposed method exhibits batch sensitivity. Excessive $B$ introduces over-regularization while Insufficient $B$ empirically degrades convergence stability.
TABLE I: Classification accuracy (%) on the CIFAR-100 validation set. ”Homogeneous Architectures” and ”Heterogeneous Architectures” represent whether teacher and student networks share the same or different architectures, respectively. $\mathrm{\Delta}$ denotes the improved performance compared to vanilla KD. All results are the average over five trials.
Method Homogeneous Architectures Heterogeneous Architectures
Teacher RN56 RN110 RN32$\times$4 W-40-2 W-40-2 VGG13 RN32$\times$4 W-40-2 VGG13 RN50 RN32$\times$4
Acc 72.34 74.31 79.42 75.61 75.61 74.64 79.42 75.61 74.64 79.34 79.42
Student RN20 RN32 RN8$\times$4 W-16-2 W-40-1 VGG8 SN-V1 SN-V1 MN-V2 MN-V2 SN-V2
Acc 69.06 71.14 72.50 73.26 71.98 70.36 70.50 70.50 64.60 64.60 71.82
Feature-based Distillation
FitNet[61] 69.21 71 06 73.50 73.58 72.24 71.02 73.59 73.73 64.14 63.16 73.54
AT[31] 70.55 72.31 73.44 74.08 72.77 71.43 71.73 73.32 59.40 58.58 72.73
RKD[56] 69.61 71.82 71.90 73.35 72.22 71.48 72.28 72.21 64.52 64.43 73.21
CRD[72] 71.16 73.48 75.51 75.48 74.14 73.94 75.11 76.05 69.73 69.11 75.65
OFD[24] 70.98 73.23 74.95 75.24 74.33 73.95 75.98 75.85 69.48 69.04 76.82
ReviewKD[7] 71.89 73.89 75.63 76.12 75.09 74.84 77.45 77.14 70.37 69.89 77.78
Logits-based Distillation
KD[25] 70.66 73.08 73.33 74.92 73.54 72.98 74.07 74.83 67.37 67.35 74.45
DML[91] 69.52 72.03 72.12 73.58 72.68 71.79 72.89 72.76 65.63 65.71 73.45
TAKD[52] 70.83 73.37 73.81 75.12 73.78 73.23 74.53 75.34 67.91 68.02 74.82
DKD[92] 71.97 74.11 76.32 76.24 74.81 74.68 76.45 76.70 69.71 70.35 77.07
CKD 72.12 74.48 76.76 76.28 75.14 74.98 77.72 77.47 70.11 70.62 78.18
$\mathrm{\Delta}$ +1.46 +1.40 +3.43 +1.36 +1.60 +2.00 +3.65 +2.64 +2.74 +3.27 +3.73
TABLE II: Top-1 and Top-5 accuracy (%) on the ImageNet-1K validation set.(a) ResNet34 serves as the teacher, while ResNet18 acts as the student. (b) ResNet50 serves as the teacher, while MobileNetV1 acts as the student. $\mathrm{\Delta}$ denotes the performance improvement compared to vanilla KD. All results are the average over three trials.
Baselines Feature-based Distillation Logits-based Distillation
Setting Metric Teacher Student AT[87] OFD[24] CRD[72] ReviewKD[7] KD[25] DKD[92] CKD $\mathrm{\Delta}$
(a) Top-1 73.31 69.75 70.69 70.81 71.17 71.61 70.66 71.70 72.24 +1.58
Top-5 91.42 89.07 90.01 89.98 90.13 90.51 89.88 90.41 90.81 +0.93
(b) Top-1 76.16 68.87 69.56 71.25 71.37 72.56 68.58 72.05 72.97 +4.39
Top-5 92.86 88.76 89.33 90.34 90.41 91.00 88.98 91.05 91.36 +2.38
TABLE III: Top-1 and Top-5 accuracy (%) on the Places365 validation set.(a) ResNet34 serves as the teacher, while ResNet18 acts as the student. (b) ResNet50 serves as the teacher, while MobileNetV1 acts as the student. $\mathrm{\Delta}$ denotes the performance improvement compared to vanilla KD. All results are the average over three trials.
Baselines Feature-based Distillation Logits-based Distillation
Setting Metric Teacher Student AT[87] OFD[24] CRD[72] ReviewKD[7] KD[25] DKD[92] CKD $\mathrm{\Delta}$
(a) Top-1 54.57 54.07 54.62 55.22 54.92 55.33 55.30 55.43 55.47 +0.17
Top-5 85.05 84.44 84.90 85.25 85.15 85.42 85.49 85.35 85.38 -0.11
(b) Top-1 55.51 51.74 52.05 55.19 54.71 54.83 53.07 55.25 55.51 +2.47
Top-5 85.51 82.29 82.67 85.16 84.86 85.13 83.48 85.50 85.41 +1.93
@@
Given the logit matrix $\boldsymbol{T}\in\mathbb{R}^{c\times n}$, and assume that there are $c$ classes, each class containing $m$ samples. The negative logit space $\boldsymbol{T}_{/i}\in\mathbb{R}^{c\times(n-1)}$ for the $i$-th sample satisfies,
@@
This rank constraint implies that the effective degrees of freedom in the negative space scale with the class count $c$ rather than the sample count $n$. Consequently, the intrinsic dimensionality of negative pairs is governed by category correlations, reducing the required sampling complexity from $\mathcal{O}(n)$ to $\mathcal{O}(c)$. Furthermore, compared to class-wise contrastive learning approaches, our method achieves efficient training through well-designed triples.
TABLE IV: Object detection performance on the MS-COCO dataset. We employ the two-stage Faster RCNN with FPN as the detector. Teacher-student pairs are ResNet101 & ResNet-18, ResNet101 & ResNet50, and ResNet50 & MobileNetV2, respectively. $\mathrm{\Delta}$ represents the improved performance compared to ReviewKD.
     Methods        RN101 & RN18        RN101 & RN50        RN50 & MN-V2
@@
TABLE V: Instance segmentation performance on the MS-COCO dataset. We take Mask R-CNN with FPN as the detector. Teacher-student pairs are ResNet101 & ResNet-18, ResNet101 & ResNet50, and ResNet50 & MobileNetV2, respectively. $\mathrm{\Delta}$ represents the improved performance compared to ReviewKD.
     Method AP AP50 AP75 APS APM APL
     RN101 38.63 60.45 41.28 19.48 41.33 55.29
     RN18 31.25 51.07 33.10 14.18 32.80 45.53
     ReviewKD [7] 33.62 53.91 35.96 15.03 35.31 50.30
     CKD+ReviewKD 34.59 55.01 36.82 15.68 36.35 51.45
     $\mathrm{\Delta}$ +0.97 +1.10 +0.86 +0.65 +1.04 +1.15
     RN101 38.63 60.45 41.28 19.48 41.33 55.29
     RN50 35.24 56.32 37.49 15.03 35.31 50.30
     ReviewKD [7] 36.98 58.13 39.60 17.54 39.57 53.19
     CKD+ReviewKD 37.89 59.45 40.61 18.40 40.62 54.91
     $\mathrm{\Delta}$ +0.91 +1.32 +1.01 +0.86 +1.05 +1.72
     RN50 35.24 56.32 37.49 15.03 35.31 50.30
     MN-V2 28.37 47.19 29.95 12.09 29.01 41.70
     ReviewKD [7] 31.56 50.70 33.44 12.76 32.44 47.39
     CKD+ReviewKD 32.80 52.64 34.96 14.67 34.38 48.84
     $\mathrm{\Delta}$ +1.24 +1.94 +1.52 +1.91 +1.94 +1.45
For classic methods with $n$ samples, a total of $\mathcal{O}(nm)$ triples should be selected for each sample,
$\binom{m-1}{1}\binom{n-m}{1}=nm-m^{2}-n+m\propto\mathcal{O}(nm)$ (23)
Eq. 23 shows that for each sample, these methods randomly choose one sample from the same class, including $m-1$ selections, and one sample from a different class, yielding a total of $\mathcal{O}(nm)$ triples. In contrast, CRD only needs to select one sample from different classes, as the positive pair is derived from the same sample,
$\binom{1}{1}\binom{n-m}{1}=n-m\propto\mathcal{O}(n).$ (24)
However, the negative space $\mathcal{O}(n)$ is extensive, requiring a memory bank of size $K$ to store historical features. There are $K$ triples available for training each sample. In contrast, the proposed method requires sampling only $\mathcal{O}(c)$ triples,
$\binom{1}{1}\binom{c}{1}=c\propto\mathcal{O}(1),$ (25)
where $c\ll n$ We randomly shuffle data samples and select $c$ non-overlapping samples sequentially. Therefore, our sample-wise formulation exhibits superior training efficiency compared to the class-wise contrastive formulation.
Finally, we show the training procedure in Algorithm 1.
IV Experiments
IV-A Image Classification
IV-A1 Datasets
We conduct experiments on three benchmark datasets, including CIFAR-100 [32], ImageNet-1K [13] and Places365 [93]. The CIFAR-100 dataset contains 100 fine-grained classes, with each class consisting of 600 images. The dataset is divided into 50,000 training images and 10,000 validation images. The ImageNet-1K dataset, a large-scale classification dataset, consists of 1000 classes, with approximately 1.28 million training images and 50,000 validation images. The Places365 dataset [93] is a large-scale scene recognition dataset containing 365 scene categories. This dataset includes 1.8 million training images and 36,500 validation images.
@@ IV-A2 Experimental Setup
We evaluate eleven isomorphic or heterogeneous architectures on CIFAR-100 and comparison methods include vanilla KD [25], AT [87], FitNet [61], RKD [56], CRD [72], OFD [24], ReviewKD [7], DML[91], TAKD[52], and DKD [92]. For the large-scale ImageNet-1K and Places365 datasets, we focus on representative architecture pairs: ResNet34 & ResNet18 and ResNet50 & MobileNetV1. We compares against vanilla KD [25], AT [87], OFD [24], CRD [72], ReviewKD [7], and DKD [92], reporting both top-1 and top-5 recognition rates across all experiments.
IV-A3 Implementation Details
All experiments are conducted on NVIDIA RTX 3090 GPUs with dataset-specific settings. For CIFAR-100, we set $\alpha$ to 100 and employ a cosine annealing learning rate scheduler. For homogeneous and heterogeneous networks, we set batch sizes to 64 and 32, with an initial learning rate of 0.05 and 0.01. For ImageNet-1K, we use cosine annealing with an initial learning rate of 0.2. We set the parameter $\alpha$ to 10 and the batch size to 512. For Places365, we set the parameter $\alpha$ to 10 and the batch size to 256.
@@ IV-A3 Implementation Details
Method Homogeneous Heterogeneous
Teacher RN56 W-40-2 W-40-2 VGG13 RN50 RN32$\times$4
@@
Student RN20 W-16-2 W-40-1 MN-V2 MN-V2 SN-V2
@@
TABLE VII: Ablation studies of triple design on the CIFAR-100 dataset. Experiments are conducted on both homogeneous and heterogeneous teacher-student pairs. Due to the fixed teacher, triples $\left(\mathbf{t}_{i},\mathbf{s}_{i},\mathbf{t}_{j}\right)$ are meaningless and excluded from this table.
Triple Homogeneous Heterogeneous
Teacher RN56 W-40-2 W-40-2 VGG13 RN50 RN32$\times$4
@@
Student RN20 W-16-2 W-40-1 MN-V2 MN-V2 SN-V2
@@
$\left({s}_{i},{t}_{i},{s}_{j}\right)$ 71.76 75.31 74.40 70.04 68.93 78.02
$\left({s}_{i},{t}_{i},{t}_{j}\right)$ 71.69 75.88 74.43 70.09 69.68 77.06
$\boldsymbol{\left({t}_{i},{s}_{i},{s}_{j}\right)}$ 72.12 76.28 75.14 70.11 70.62 78.18
@@ IV-A4 Performance on CIFAR-100
We conduct classification experiments by using eleven combinations of isomorphic or heterogeneous teacher-student networks, including ResNet [22], WideResNet [75], VGG [65], ShuffleNet [90, 47], and MobileNet [62].
Table I shows classification performance using both homogeneous and heterogeneous teacher-student architectures. Our method achieves competitive performance against feature-based and logit-based methods. The proposed CKD achieves an average improvement of 2.20% over vanilla KD across all 11 settings, with a maximum gain of 3.73%. Our method surpasses state-of-the-art feature distillation approaches, with an average accuracy improvement of 1.00% against CKD. Moreover, CKD shows stronger improvements in heterogeneous settings compared to homogeneous configurations. The superior performance of sample-wise contrastive learning over class-wise alternatives is evident in heterogeneous scenarios. The sample-wise contrastive mechanism appears particularly effective when dealing with heterogeneous architectures. Unlike feature-based methods requiring explicit layer alignment, CKD creates architecture-agnostic embeddings with logit alignment, better bridging the representation gap between different networks.